# Social Genome: Grounded Social Reasoning Abilities of Multimodal Models

### v2

@@ 1 Introduction
Developing algorithms for multimodal social reasoning will be essential to advance artificial intelligence (AI) systems with social intelligence Mathur et al. (2024). When AI systems reason about human social interactions, it is important for systems to have the ability to generate interpretable explanations with accurate, grounded references to fine-grained multimodal behaviors and external social knowledge concepts that inform inferences. Figure 1 visualizes these aspects of multimodal social reasoning. This capability is especially important for AI systems reasoning about human social interactions in high-stakes domains, such as healthcare agents and assistive robots.
@@
Progress towards improving multimodal social reasoning abilities of models has been limited by a lack of evaluation tasks – measuring a capability is an essential first step towards advancing it. We introduce Social Genome, the first benchmark for grounded multimodal social reasoning that includes 272 videos of face-to-face interactions and 1,486 human-annotated reasoning traces explaining inferences about social information in these videos. Across these traces, Social Genome contains 5,777 social reasoning steps. Each reasoning step is tagged with the modality of information being referenced: visual, verbal, and vocal cues from social interactions in videos, and external knowledge of contextual information that human annotators used to perform social inferences (information external to stimuli in videos). Reasoning traces in Social Genome reference over 11,000 entities (people, objects, concepts), over 5000 multimodal cues, and over 2,900 external knowledge observations. Social Genome is the first social reasoning benchmark to include external knowledge and dense reasoning traces. A sample human-annotated reasoning trace is provided in Figure 2.
This paper defines metrics to holistically assess semantic and structural aspects of model-generated social reasoning traces. We demonstrate the utility of Social Genome by using these metrics to distill insights regarding social reasoning capabilities and limitations in state-of-the-art (SOTA) multimodal models. For example, we find that models struggle to perform well under both zero-shot and in-context learning (ICL) settings, demonstrating the significant challenge of building this understudied form of reasoning in models. Our findings contribute novel insights regarding gaps and opportunities for future research to improve the grounded social reasoning abilities of multimodal models.
@@ 2 Background
The ability of models to perform social reasoning on multimodal human interactions, in particular face–to–face, embodied, real-world social interactions, has been comparatively understudied. Key benchmarks include the video QA tasks of Social-IQ 1.0 Zadeh et al. (2019) and Social-IQ 2.0 Wilf et al. (2023); both examine model QA accuracy when answering questions about social interactions in videos. SOTA models have struggled to perform well on Social-IQ 2.0 Xie and Park (2023); Pirhadi et al. (2023); Li et al. (2024c); Agrawal et al. (2024); Chen et al. (2024). The community’s prior focus on QA accuracy to assess social reasoning ability has not enabled researchers to study the extent to which models can effectively reference multimodal cues and external knowledge informing inferences. Models with high accuracy on QA tasks can perform poorly at generating valid or comprehensive reasoning traces Jhamtani and Clark (2020); Gu et al. (2023). Model reasoning trace quality has not been studied before in the context of multimodal social reasoning. We introduce Social Genome as the first benchmark to study grounded, fine-grained social reasoning abilities of multimodal models.
@@ 3.1 Sourcing Seed Videos and Questions
The Social Genome benchmark contains 272 seed videos and 1486 questions adapted from the publicly-available Social-IQ 2.0 dataset Wilf et al. (2023) (details on data sourcing in Appendix A.1). Videos include real-world face-to-face dyadic and multi-party interactions (1 minute per video, $\sim$4.5 total hours), and questions probe social dynamics, behaviors, emotions, and cognitive states of both individuals and groups. From seed videos and questions, Social Genome introduces a new set of 1486 human reasoning traces with 5700+ steps that answer these questions.
@@ 3.2 Task Notation
Given a video $V$ depicting a social interaction, a question $Q$ about social content in the interaction, and a corresponding set of answer options $A=\{A_{\text{correct}},A_{\text{incorrect}_{1}},A_{\text{incorrect}_{2}},A_{\text{incorrect}_{3}}\}$, a model performing the Social Genome task must generate a reasoning trace $R=\{e_{1},e_{2},\dots,e_{n}\}$, where each reasoning step $e_{i}$ represents a single piece of evidence contributing toward the social inference to select an answer $A_{\text{a}}$ from $A$. Each reasoning step $e_{i}$ must be tagged with two attributes: (1) a modality tag $m_{i}\in\{\textit{visual, verbal, vocal, }n/a\}$ indicating the communication modality of the evidence and (2) an external knowledge tag $k_{i}\in\{\text{{yes}, {no}}\}$, indicating whether the evidence references external knowledge of contextual information. This task to generate $\mathit{R}$ and answer question Q evaluates a model’s ability to extract and reference multimodal aspects of human communication and knowledge informing social inferences. Given the input tuple $(V,Q,A)$, each model performing the Social Genome task will produce an output tuple $(A_{\text{a}},R)$. Metrics in Social Genome study the social inference accuracy of $A_{a}$ and the semantic and structural aspects of social reasoning in $R$.
@@ 3.3 Social Reasoning Trace Annotations
Humans perceive and build upon low-level observations of fine-grained behaviors (e.g, shifts in body language) and high-level, top-down processing (e.g., implicit situational knowledge) when interpreting social scenes Baird and Baldwin (2001); Bodenhausen and Morales (2013). Annotators were instructed to reference any low-level and high-level evidence they relied upon to answer questions: for example, low-level evidence might be "the woman takes a step back with her mouth wide open (visual cue)" and high-level evidence that interprets that low-level cue might be "the woman is surprised (external knowledge regarding how ’surprise’ might manifest"). For each step, annotators tagged the perceptual modality referenced (visual, verbal, vocal, n/a).
@@
Annotators tagged each reasoning step with $yes$ or $no$ to indicate whether external knowledge was referenced. External knowledge includes contextual norms, cultural expectations, and prior understanding of social commonsense Forguson and Gopnik (1988) that goes beyond stimuli in the video. For example, if a man raises his arm and the annotator recognizes his movement as a "high five," the identification of that gesture and its interpretation as "triumphant" is based on external knowledge of social norms.
@@
Trained experts validated each Prolific annotation. They watched each video, read each QA tuple, and read the annotation to ensure that traces represented valid reasoning, had correct modality and external knowledge tags, and referred to relevant information. Cases of incomplete annotation or deviation from instructions were fixed, as discussed in Appendix A.3.
@@ 3.4 Dataset Statistics
Social Genome contains 1486 detailed human-annotated social reasoning traces with 5,777 total steps, 3.89 $\pm$ 1.68 steps per trace (minimum of 1 step and maximum of 10 steps), 43 $\pm$ 26 words per trace, and 11$\pm$ 5 words per step. Reasoning steps draw on multimodal evidence, with 44% of steps referencing visual cues, 27% referencing verbal cues, and 17% referencing vocal cues. Overall, 77% of traces referenced at least one visual cue, 63% referenced at least one verbal cue, and 47% referenced at least one vocal cue.
External knowledge plays a critical role in Social Genome: 51% of reasoning steps referenced external knowledge, with each trace referencing an average of 2 pieces of external knowledge. With spaCy named entity recognition (NER) Honnibal et al. (2020), we found 11,253 entities (people, objects, concepts) mentioned, with 7.6 unique entities and 2.23 emotions referenced per reasoning trace, demonstrating the high density of annotations.
@@ 3.5 Social Reasoning Metrics and Statistics
We develop a set of metrics to holistically evaluate semantic and structural aspects of social reasoning traces generated by models performing tasks in the Social Genome benchmark. Collectively, these metrics reveal strengths and weaknesses in model social reasoning and multimodal grounding abilities and the extent to which model traces differ from human reasoning. For each sample, we compute the following metrics between model reasoning trace $\mathit{R_{M}}$ with $n$ steps $e_{1}...e_{n}$ and the corresponding human trace $\mathit{R_{H}}$ with $m$ steps $h_{1}...h_{m}$.
@@
Measures the number of reasoning steps in $\mathit{R_{M}}$ with a semantic similarity above a threshold $\tau$, when compared to any step in $\mathit{R_{H}}$. This metric is computed as:
@@
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with visual evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$, with respect to the presence of visual evidence (max value is number of visual steps in $\mathit{R_{H}}$).
@@
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with verbal evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$, with respect to the presence of verbal evidence (max value is the number of verbal steps in $\mathit{R_{H}}$).
@@
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with vocal evidence to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$, with respect to the presence of vocal evidence (max value is the number of vocal steps in $\mathit{R_{H}}$).
@@
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with external knowledge evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$, with respect to external knowledge references (max value is the number of external knowledge steps in $\mathit{R_{H}}$).
@@
Figure 3: Performance of models across number of few-shot ICL samples ($k$) and ground truth noted in gray.
With these metrics, Social Genome enables a holistic study of multimodal, grounded social reasoning.
The first six metrics focus on social inference accuracy, semantic similarity, and structural similarity between model and human reasoning traces. The final six metrics focus on fine-grained multimodal evidence and external knowledge referenced by models, in comparison to evidence referenced by humans.
@@ 4.1 Quantitative Results and Insights
Figure 3 visualizes each model’s average performance for our 12 metrics [^1] Visualized metrics are normalized to [0,1] by human baselines, except for those with an upper bound of 1 by definition (Accuracy, Similarity-Trace, Similarity-Step, DifferenceSequence) and absolute measures (NumSteps)., with human baselines in gray. Results tables (Tables 3, 4, 5) are in Appendix C. Key findings are discussed below.
@@
Human social inference ability is substantially higher than all models, as seen in results from the Accuracy metric. Gemini-1.5-Flash and GPT-4o achieve the highest accuracies of 74.4% and 71.0% respectively ($k$ = 0), approximately 10-15% lower than human annotator accuracy (85.3%), answering $\sim$30of questions incorrectly. Closed-source models outperform open-source models in social inference. The highest-performing open-source model was LLaVA-Video at 62.9% ($k$ = 0). Gemini-1.5-Flash and GPT-4o are much larger than open-source models, suggesting that increased model scale is useful, but not sufficient, for accurate social inference. Video-ChatGPT and VideoChat2 perform 30-40% lower than other open-source models across all values of $k$. These models have the smallest context length and process the fewest frames, constraints which may influence performance.
Social inference accuracy for models decreased as the number of few-shot ICL samples increased, with the exception of GPT-4o, which demonstrated a slight improvement at $k$ = 16. Few-shot ICL conditions language models on tasks by providing examples of inputs and outputs Liu et al. (2024) and can be viewed as a form of inductive reasoning, as the model is tasked with inferring generalizable rules from a set of examples. This technique has improved model reasoning abilities in domains such as mathematics and code generation Dong et al. (2024); Zhou et al. (2022); Patel et al. (2023), which have explicit rules and formal structure Galotti (1989). In contrast to these domains, social reasoning often operates with implicit rules, less formal structure Perkins (1989) and ambiguity in premises Mathur et al. (2024). Our findings suggest that few-shot ICL may not be an effective approach to elicit multimodal social reasoning abilities. Additional experiments (provided in Appendix D) demonstrate that chain-of-thought prompting did not improve model accuracy, and models struggled to perform inferences relying on implicit and contextual knowledge. Our finding from Social Genome that few-shot ICL is insufficient to elicit multimodal social reasoning aligns with findings from unimodal experiments on Social-IQa and ToMi datasets Le et al. (2019); Sap et al. (2019, 2022); Kim et al. (2023).
@@ 4.2 Human and Qualitative Evaluation
We conducted human evaluation of model reasoning traces. Trained annotators analyzed 48 samples from all models, with model names and $k$ values anonymized. Traces were rated to assess references to low-level cues (fine-grained), information cross-referenced across steps (compositional), relevant evidence (comprehensive), correctness of modality tags, and validity of reasoning (criteria are in Appendix E). Annotator agreement was computed with Cohen’s Kappa $\kappa$ Cohen (1960).
Reasoning traces from Gemini-1.5-Flash and GPT-4o received the highest ratings for fine-grained ($\kappa$ = 0.87), compositional ($\kappa$ = 0.92), comprehensive ($\kappa$ = 0.94), and valid ($\kappa$ = 0.94) reasoning, followed by LLaVA-Video and LongVA, with VideoChatGPT and VideoChat2 rated lowest. These findings validate model performance trends observed in Section 4.1. Modality tag correctness across samples was 98%, supporting the validity of the Social Genome automated modality metrics and reasoning trace processing (Appendix B.4).
We visualize representative traces in Figure 4 from a human annotator, Gemini-1.5-Flash, and LLaVA-Video (closed-source and open-source models with highest inference accuracy at $k$ = 0).
@@
This figure illustrates the strong ability of Gemini-1.5-Flash to reference and integrate multimodal cues and external knowledge during social reasoning. In contrast, LLaVA-Video did not reference low-level cues and based its reasoning upon an incorrect premise, and this error propagated to an incorrect inference. In addition, the human trace referred to several fine-grained, subtle behaviors (e.g., lip movements) that are not present in the model traces, yet do influence the scene interpretation. Similarity metrics and modality-specific metrics in Section 3.5 serve as automated proxies to estimate these types of information gaps between human and model reasoning, as shown in these qualitative examples. Our findings motivate future work on model training data, training approaches, and architectures that better capture fine-grained cues, as well as algorithms to handle error propagation in social reasoning.
@@
Figure 4 qualitatively illustrates the hierarchical structure of human social reasoning traces, in which low-level cues (e.g., brief lip movement) are referenced, combined, and re-interpreted as intermediate evidence for further reasoning. This “forking" reasoning structure is common for humans interpreting everyday situations, unlike the linear “long chain" structure for formal reasoning in domains such as mathematics Perkins (1989); Galotti (1989). As seen in Figure 4, compared to human traces, model traces have flatter structures with less hierarchical composition of social cues. Flatter structures have the potential to treat cues as isolated evidence and overlook intermediate evidence in reasoning paths. Our findings motivate future work to train models capable of more hierarchical social reasoning.

### v3

@@ 1 Introduction
Developing algorithms for multimodal social reasoning will be essential to advance artificial intelligence (AI) systems with social intelligence Mathur et al. (2024). When AI systems reason about human social interactions, it is important for systems to have the ability to generate explanations with accurate, grounded references to fine-grained multimodal behaviors and external knowledge concepts informing inferences. Figure 1 visualizes these aspects of multimodal social reasoning. This capability is especially important for AI systems reasoning about interactions in high-stakes domains, such as healthcare and assistive robots.
@@
Progress towards improving multimodal social reasoning in models has been limited by a lack of evaluation tasks – measuring a capability is an essential first step towards advancing it. To address this challenge, we introduce Social Genome, the first benchmark for grounded multimodal social reasoning that includes 272 videos of face-to-face interactions and 1,486 human-annotated reasoning traces explaining inferences about social information in these videos. Across these traces, Social Genome contains 5,777 social reasoning steps. Each reasoning step is tagged with the modality of information being referenced: visual, verbal, and vocal cues from social interactions in videos, and external knowledge of contextual information that human annotators used to perform social inferences (information external to stimuli in videos). Reasoning traces in Social Genome are dense with references to over 11,000 entities (people, objects, concepts), over 5000 multimodal cues, and over 2,900 external knowledge observations. Social Genome is the first social reasoning benchmark111cmu-multicomp-lab.github.io/social-genome that includes external knowledge and dense reasoning traces. A sample human-annotated reasoning trace is visualized in Figure 2.
Benchmark Social Reasoning Task Focus Reasoning Traces Fine-Grained Evaluation External Knowledge
Social Genome (this paper)
Social-IQ 1.0 Zadeh et al. (2019)
Social-IQ 2.0 Wilf et al. (2023)
TVQA Lei et al. (2018)
MovieQA Tapaswi et al. (2016)
MEmoR Shen et al. (2020)
Ego4D-Social Grauman et al. (2022)
Table 1: Social Genome enables the study of fine-grained, grounded social reasoning in multimodal models.  = provided;  = not provided. Ego4D has a social signal perception task, which differs from a social reasoning task.
This paper defines metrics to holistically assess semantic and structural aspects of model-generated social reasoning traces. We demonstrate the utility of Social Genome by using these metrics to distill insights regarding social reasoning capabilities and limitations in state-of-the-art (SOTA) models. For example, we find that models struggle to perform well under both zero-shot and in-context learning (ICL) settings, demonstrating the significant challenge of building this understudied form of reasoning in models. Our findings contribute novel insights regarding gaps and opportunities for future research to improve grounded social reasoning abilities of multimodal models.
@@ 2 Background
The ability of models to reason about multimodal social interactions, in particular face–to–face, embodied, real-world social interactions, has been comparatively understudied. Key benchmarks include the video QA tasks of Social-IQ 1.0 Zadeh et al. (2019) and Social-IQ 2.0 Wilf et al. (2023); both examine model QA accuracy when answering questions about social interactions in videos. SOTA models have struggled to perform well on Social-IQ 2.0 Xie and Park (2023); Pirhadi et al. (2023); Li et al. (2024c); Agrawal et al. (2024); Chen et al. (2024). Prior focus on QA accuracy to assess social reasoning ability has not enabled researchers to study the extent to which models can effectively reference fine-grained multimodal cues and external knowledge informing inferences. Models with high accuracy on QA tasks can perform poorly at generating valid or comprehensive reasoning traces Jhamtani and Clark (2020); Gu et al. (2023), motivating the creation of benchmarks with fine-grained evaluation that goes beyond QA tasks. We introduce Social Genome as the first benchmark to study grounded, fine-grained social reasoning in multimodal models. Table 1 summarizes the novelty of Social Genome  relative to prior human-centered video understanding tasks.
@@ 3.1 Sourcing Seed Videos and Questions
Social Genome contains 272 seed videos and 1486 questions adapted from the Social-IQ 2.0 dataset Wilf et al. (2023) (details in Appendix A.1). Videos include real-world face-to-face interactions (1 min per video, $\sim$4.5 hours); questions probe behaviors, emotions, and cognitive states of individuals and groups. Social Genome introduces a new set of 1486 human reasoning traces with 5700+ steps that answer these questions.
@@ 3.2 Task Notation
Given a video $V$, a question $Q$ about social interactions in the video, and answer options $A=\{A_{\text{correct}},A_{\text{incorrect}_{1}},A_{\text{incorrect}_{2}},A_{\text{incorrect}_{3}}\}$, a model performing the Social Genome task must generate a reasoning trace $R=\{e_{1},e_{2},\dots,e_{n}\}$, where each reasoning step $e_{i}$ represents a single piece of evidence contributing toward the social inference to select an answer $A_{\text{a}}$ from $A$. Each reasoning step $e_{i}$ must be tagged with two attributes: (1) a modality tag $m_{i}\in\{\textit{visual, verbal, vocal, }n/a\}$ indicating the communication modality of the evidence and (2) an external knowledge tag $k_{i}\in\{\text{{yes}, {no}}\}$, indicating whether the evidence references external knowledge of contextual information. This task to generate $\mathit{R}$ and answer question Q evaluates a model’s ability to extract and reference multimodal aspects of human communication and knowledge informing social inferences. Given the input tuple $(V,Q,A)$, each model performing the Social Genome task will produce an output tuple $(A_{\text{a}},R)$. Metrics in Social Genome study the social inference accuracy of $A_{a}$ and the semantic and structural aspects of social reasoning in $R$.
@@ 3.3 Social Reasoning Trace Annotations
Humans build upon low-level observations of fine-grained behaviors (e.g, shifts in body language) and high-level, top-down processing (e.g., implicit situational knowledge) when interpreting social scenes Baird and Baldwin (2001); Bodenhausen and Morales (2013). Annotators were instructed to reference any low-level and high-level evidence they relied upon to answer questions: for example, low-level evidence might be "the woman takes a step back with her mouth wide open (visual cue)" and high-level evidence that interprets that low-level cue might be "the woman is surprised (external knowledge regarding how ’surprise’ might manifest"). For each step, annotators tagged the modality referenced (visual, verbal, vocal, n/a).
@@
Annotators tagged each reasoning step with $yes$ or $no$ to indicate whether external knowledge was referenced. External knowledge includes contextual norms, cultural expectations, and prior understanding of social commonsense Forguson and Gopnik (1988) that goes beyond stimuli in the video. For example, if a man raises his arm and the annotator recognizes his movement as a "high five," the identification of the gesture is based on external knowledge.
@@
Trained experts validated each Prolific annotation. They watched each video, read each QA tuple, and read the annotation to ensure that traces represented valid reasoning, had correct modality and external knowledge tags, and referred to relevant information. Cases of incomplete annotation or deviation from instructions were fixed (details in Appendix A.4).
@@ 3.4 Dataset Statistics
Social Genome contains 1486 human-annotated reasoning traces with 5,777 total steps, 3.89 $\pm$ 1.68 steps per trace (min of 1 step and max of 10 steps), 43 $\pm$ 26 words per trace, and 11$\pm$ 5 words per step. Reasoning steps draw on multimodal evidence; 44% of steps reference visual cues, 27% reference verbal cues, and 17% reference vocal cues. Overall, 77% of traces reference at least one visual cue, 63% reference at least one verbal cue, and 47% reference at least one vocal cue.
External knowledge plays a critical role in Social Genome: 51% of reasoning steps referenced external knowledge, with each trace referencing an average of 2 pieces of external knowledge. With spaCy named entity recognition (NER) Honnibal et al. (2020), we found 11,253 entities (people, objects, concepts) mentioned, with 7.6 unique entities and 2.23 emotions referenced per reasoning trace, demonstrating the high density of annotations. Additional details regarding dataset statistics are in Section 7 and Appendix A.2.
@@ 3.5 Social Reasoning Metrics and Statistics
We develop metrics to evaluate semantic and structural aspects of social reasoning traces generated by models performing tasks in the Social Genome benchmark. Collectively, these metrics reveal strengths and weaknesses in model social reasoning and multimodal grounding abilities and the extent to which model traces differ from human reasoning. This multi-dimensional evaluation mitigates against models hacking individual metrics to achieve higher scores. For each sample, we compute the following metrics between model reasoning trace $\mathit{R_{M}}$ with $n$ steps $e_{1}...e_{n}$ and the corresponding human trace $\mathit{R_{H}}$ with $m$ steps $h_{1}...h_{m}$.
@@
Measures the number of steps in $\mathit{R_{M}}$ with a similarity above threshold $\tau$, when compared to any step in $\mathit{R_{H}}$:
@@
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with visual evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$ (max value is number of visual steps in $\mathit{R_{H}}$).
@@
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with verbal evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$ (max value is the number of verbal steps in $\mathit{R_{H}}$).
@@
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with vocal evidence to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$ (max value is the number of vocal steps in $\mathit{R_{H}}$).
@@
Measures the number of steps in both $\mathit{R_{M}}$ and $\mathit{R_{H}}$ with external knowledge evidence, to evaluate how closely $\mathit{R_{M}}$ aligns with $\mathit{R_{H}}$ (max value is the number of external knowledge steps in $\mathit{R_{H}}$).
@@
Figure 3: Performance of models across number of few-shot ICL samples ($k$) and ground truth noted in gray. The first six metrics focus on social inference accuracy, semantic similarity, and structural similarity between model and human reasoning traces. The final six metrics focus on fine-grained multimodal evidence and external knowledge referenced by models, in comparison to evidence referenced by humans. With these metrics, Social Genome enables a holistic study of multimodal, grounded social reasoning.
@@ 4.1 Quantitative Results and Insights
Figure 3 visualizes each model’s average performance for our 12 metrics [^2] Visualized metrics are normalized to [0,1] by human baselines, except for those with an upper bound of 1 by definition (Accuracy, Similarity-Trace, Similarity-Step, DifferenceSequence) and absolute measures (NumSteps)., with human baselines in gray. Results tables (Tables 4, 5, 6) are in Appendix C. Key findings are discussed below.
@@
Human social inference ability is substantially higher than all models, as seen in results from the Accuracy metric. Gemini-1.5-Flash and GPT-4o achieve the highest accuracies of 74.4% and 71.0% respectively ($k$ = 0), approximately 10-15% lower than human annotator accuracy (85.3%), answering $\sim$30of questions incorrectly. Closed-source models outperform open-source models in social inference. The highest-performing open-source model was LLaVA-Video at 62.9% ($k$ = 0). Gemini-1.5-Flash and GPT-4o are much larger than open-source models, suggesting that increased model scale is useful, but not sufficient, for social inference. Video-ChatGPT and VideoChat2 perform 30-40% lower than other open-source models across all values of $k$. These models have the smallest context length, constraints which may influence performance.
Social inference accuracy for models decreased as the number of few-shot ICL samples increased, with the exception of GPT-4o, which demonstrated a slight improvement at $k$ = 16. Few-shot ICL conditions language models on tasks by providing examples of inputs and outputs Liu et al. (2024) and can be viewed as a form of inductive reasoning, as the model is tasked with inferring generalizable rules from a set of examples. This technique has improved model reasoning abilities in domains such as mathematics and code generation Dong et al. (2024); Zhou et al. (2022); Patel et al. (2023), which have explicit rules and formal structure Galotti (1989). In contrast to these domains, social reasoning often operates with implicit rules, less formal structure Perkins (1989) and ambiguity in premises Mathur et al. (2024). Our findings suggest that few-shot ICL may not be an effective approach to elicit multimodal social reasoning abilities. Additional experiments using Social Genome samples as a form of supervision for models (discussed in Appendix D) demonstrate that chain-of-thought prompting did not improve model accuracy, and models struggled to perform inferences relying on implicit and contextual knowledge. Our finding that few-shot ICL is insufficient to elicit multimodal social reasoning aligns with findings from unimodal experiments on Social-IQa and ToMi datasets Le et al. (2019); Sap et al. (2019, 2022); Kim et al. (2023).
@@ 4.2 Human and Qualitative Evaluation
We conducted human evaluation of model reasoning traces. Trained annotators analyzed 48 samples from all models, with model names and $k$ values anonymized. Traces were rated to assess references to low-level cues (fine-grained), information cross-referenced across steps (compositional), relevant evidence (comprehensive), correctness of modality tags, and validity of reasoning (details in Appendix E). Annotator agreement was computed with Cohen’s Kappa $\kappa$ Cohen (1960).
Reasoning traces from Gemini-1.5-Flash and GPT-4o received the highest ratings for fine-grained ($\kappa$ = 0.87), compositional ($\kappa$ = 0.92), comprehensive ($\kappa$ = 0.94), and valid ($\kappa$ = 0.94) reasoning, followed by LLaVA-Video and LongVA, with VideoChatGPT and VideoChat2 rated lowest. These findings validate model performance trends in Section 4.1. Modality tag correctness across samples was 98%, and metrics correlated with human judgements ($R^{2}>$ 0.75 for Gemini-1.5-Flash semantic metrics, details in Appendix E), supporting the validity of benchmark processing.
@@
Figure 4 illustrates the strong ability of Gemini-1.5-Flash to reference and integrate multimodal cues and external knowledge. In contrast, LLaVA-Video did not reference low-level cues and based its reasoning upon an incorrect premise which led to an incorrect inference. The human trace referred to fine-grained behaviors (e.g., lip movements) that are not present in the model traces, yet do influence the scene interpretation. Metrics in Section 3.5 serve as proxies to estimate these types of information gaps between human and model reasoning. Our findings motivate future work on model training and architectures that better capture fine-grained cues and handle error propagation in reasoning.
@@
Figure 4 shows the hierarchical structure of human social reasoning traces, in which low-level cues (e.g., brief lip movement) are referenced, combined, and re-interpreted as intermediate evidence for further reasoning. “Forking" reasoning structures are common for humans interpreting everyday situations, unlike the linear “long chain" structures for formal reasoning in domains such as mathematics Perkins (1989); Galotti (1989). Compared to human traces, model traces show flatter structures which have the potential to overlook intermediate evidence. Our findings motivate future work to train models capable of more hierarchical social reasoning.